Circular No. 5866

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOTICE TO SUBSCRIBERS Because of sharply increased dissemination costs, it will become neces-

sary, starting Oct. 1, to increase the subscription rates to the printed edition of these Circulars. The new monthly rates will be $12.50 at the regular (invoiced) rate and $7.50 at the special (non-invoiced) rate. At the same time, however, we are prepared to drop the requirement that subscribers to the electronic version must also subscribe to the printed edition. If subscribers notify us that they no longer require the Circulars in printed form, they can therefore utilize the Computer Service and/or receive the Circulars by e-mail for a monthly charge of only $12.50 (regular rate) or $7.50 (special rate). Furthermore, the Minor Planet Center is about to initiate a series of Minor Planet Electronic Circulars (MPECs), containing urgent information about unusual minor planets. These will also be available in the Computer Service and sent, at least initially, to all e-mail subscribers at no extra charge. They will contain (and much more extensively) most of the information about near-earth objects that has in the past been contained in the IAU Circulars (as well as some less urgent information about comets). The MPECs will not be issued in printed form, although all the relevant material in them will also be published in the monthly Minor Planet Circulars. All subscribers should be aware that this arrangement will be economically viable only if there is a significant increase in the number of subscribers to the Computer Service/e-mail delivery. If you are reading a ‘pirated’ version of this Circular, you are therefore urged to open your own subscription by contacting the postal or e-mail addresses above.

COMET McNAUGHT-TRITTON (1978 XXVII) On Aug. 2 R. H. McNaught, Anglo-Australian Observatory, informed

us that he had been able to find another image of the 1978 May 1 comet he reported on IAUC 5471. This was on a U.K. Schmidt exposure on 1978 Apr. 12 by K. P. Tritton and H. T. MacGillivray. The undersigned then suggested identity with the other comet mentioned on IAUC 5471, i.e., the one discovered by S. B. Tritton on a 1979 Mar. 5 plate (taken by W. J. Zealey) and the measurements of which were published with an erroneous date on IAUC 3662. A parabolic orbit solution (T = 1978 Aug. 25, q = 6.28 AU, i = 153^{o}) then allowed S. B. Tritton to find a faint image of the comet on a plate taken by J. A. Dawe on 1980 Jan. 23. The 1978 May 1 image was also independently noticed by D. F. Malin.

1993 September 18 Brian G. Marsden